VRT-SPECTROSCOPY IN THE TRANSLATIONAL AND LIBRATIONAL BAND REGION OF LIQUID WATER: HYDROGEN BOND TUNNELING DYNAMICS IN WATER CLUSTERS

\underline{F. N. KEUTSCH}, R. S. FELLERS, P. B. PETERSEN, M. R. VIANT, M. G. BROWN, and R. J. SAYKALLY, Department of Chemistry, University of California Berkeley, Berkeley, CA 94720.

We report the observation of a new vibration-rotation-tunneling (VRT) band of (D_{2}O)_{3} at 142.8 cm^{−1} and a set of four bands of (H_{2}O)_{3} around 520 cm^{−1}. These new bands represent the first observation of a translational and librational vibration for a water cluster. The observed VRT spectrum of (D_{2}O)_{3} at 142.8 cm^{−1}, in the translational band of the liquid, is assigned to a combination band or mixed level of the asymmetric hydrogen bond stretch and a torsional vibration. The predicted frequencies of the hydrogen bond stretching modes are too high, presumably because calculations fail to include the necessary coupling between stretching and torsional motions ^{a}. The bands of (H_{2}O)_{3} around 520 cm^{−1} lie in the librational band region of liquid water and are tentatively assigned to the out of plane librational vibration. The observation of at least four bands within 8 cm^{−1} is explained by a dramatically increased splitting of the excited state rovibrational levels by bifurcation-tunneling. The experimental results presented should therefore allow for the first exact determination of the height of the bifurcation tunneling barrier. The tunneling time scale is estimated at 2-4ps, similar to those of several important dynamical processes in bulk water ^{b}.

aW. Klopper, M. Schutz, H. P. Luthi, S. Leutwyler, J. Chem. Phys. 103, 1085 (1995).

bA. Luzar, D. Chandler, Nature 379, 55 (1996).